# Towards Precise Scaling Laws for Video Diffusion Transformers

### v1

@@ 2 Setup
  • $C_{\text{token}}=\frac{3}{4}N\left(7+\frac{n_{\text{ctx}}}{d}\right)$: The compute cost per token during training, where $d$ represents the model dimension.
@@
3 Scaling Laws for Hyperparameters
@@
These analyses underscore the importance of a refined approach to hyperparameter selection in scaling law research. To address this, we propose the precise Scaling Law for video DiT. In this section, we derive a power-law relationship linking hyperparameters to model size and data size, offering a theory-driven improvement over prior methods. This serves as the basis for our scaling law in Section 4.
@@ 3.1 Impact of model size on hyperparameters
To facilitate convergence analysis, we assume that, conditioned on the filtration $\mathcal{G}_{k}^{B}$,
$\mathcal{G}_{k}^{B}=\left\{\theta_{k},\left\{\xi_{k-1}^{(b)}\right\}_{b=1}^{B},\theta_{k-1},\left\{\xi_{k-2}^{(b)}\right\}_{b=1}^{B},\cdots,\theta_{0}\right\}$ (2)
the mini-batch gradient $g_{k}$ is an unbiased estimate on $\nabla L(\theta_{k})=G(\theta_{k})$, and the variance is bounded by $\sigma_{B}^{2}$:
@@
To simplify the analysis, we assume that $L(\theta_{k})$ is $L$-smooth for any iteration k, and we derive the convergence properties under the framework of Mini-Batch Stochastic Gradient Descent (SGD) [37]. For effective updates, the learning rate $\eta$ must satisfy $\eta\leq 1/L$. Under these conditions, the convergence of $L(\theta_{k})$ is characterized as follows (proof provided in Appendix B):
@@
As the model size increases, the complexity of the model also grows, leading to a corresponding increase in the Lipschitz constant $L$. To ensure that the upper bound on the right-hand side of Equation 5 remains manageable, both the learning rate and batch size must be adjusted with respect to $L$. Specifically, a larger $L$ necessitates a smaller learning rate $\eta$ and a larger batch size $B$ for convergence. This relationship illustrates how the scaling of model size impacts the optimal choice of hyperparameters. Unlike prior work [1], which treats model size implicitly, our analysis explicitly incorporates model size as an independent variable. This enables a more accurate estimation of optimal hyperparameters by addressing the interplay between model complexity, learning rate, and batch size in a principled manner.
@@ 3.2 Proposed $B_{\text{opt}}(N,T)$ equation
$\mathbb{D}_{\xi_{k}\sim\rho}[g_{k}]=\frac{1}{B}\Sigma_{k}\vspace{-0.5em}$ (6)
@@ 3.4 Results of optimal hyperparameters
4 Scaling Laws in Video DiT
@@ 4.1 Scaling law for optimal model size
In this section, we explore how to optimally allocate model size $N$ and training tokens $T$ within a given compute budget $C$ to maximize model performance. Building on the approach proposed in [1], we leverage the optimal hyperparameters defined in Equations 14 and 11 for our experiments. For each compute budget $[3e17,6e17,1e18,3e18,6e18]$, we conducted five experiments to ensure reliable results. As shown in Figure 6(a), we fit a parabola to each compute budget’s results to identify the minimum loss point, which corresponds to empirical optimal model size under that specific budget.
@@ 4.3.1 More parameter-efficient model sizes
We first validate our fitting results by extrapolating the compute budget to $5.85\times 10^{20}$. Under this condition, the optimal model size predicted by Equation 15 is 0.64B. Due to discrete layer selection, we conducted an experimental validation using a $14$-layer model with 719.3M parameters. According to equations Equations 11 and 14, we set the batch size to $832$ and the learning rate to $1.6\times 10^{-4}$. The resulting loss closely aligned with the prediction from Equation 16 (see Figure 5(b)), further validating our conclusions.
To validate the optimal hyperparameters yield more precise fitting results, we conducted experiments with fixed suboptimal hyperparameters to fit the relationship between optimal model size and compute budget (Appendix C). As shown in Figure 6, using fixed suboptimal hyperparameters overestimates the optimal model size, and this discrepancy grows with increasing compute budgets. At a compute budget of $1\times 10^{10}$ TFlops, the scaling law based on optimal hyperparameters predicts a model that saves $39.9\%$ of the parameters while achieving a similar validation loss. This occurs because, when computational resources are abundant relative to data, scaling the model size near the optimal size has a minimal impact on validation loss. Therefore, our scaling law precisely predict the optimal model size under optimal hyperparameters.
@@ 4.3.2 More precise validation loss predictions
The corresponding results under fixed suboptimal hyperparameters are provided in  Appendix C. We focus on the deviation in the exponent of $C$ between the predicted optimal model size and the empirical optimal model size relationships. As depicted in Figure 7, under optimal hyperparameters, the deviation results in an absolute error of $0.0148$, corresponding to a $3.57\%$ slope deviation. In contrast, as shown in Equations 39 and 38, the deviation under fixed suboptimal hyperparameters is significantly larger, with an absolute error of $0.1581$, and $30.26\%$ difference. This stark difference underscores the importance of using optimal hyperparameters to improve the precision of predicted relationships. Therefore, our scaling laws precisely predict the validation loss under optimal hyperparameters.
@@ A.3 Main Results
Based on the optimal learning rate, we fit the validation loss for any model size and training tokens (in Table 5).
@@ B.3 Stepwise Loss of Mini-Batch SGD
This estimated gradient is unbiased, and its variance decreases inversely with the batch size $B$ (see in Section 3.1 ).
@@ Appendix C Experimental Conclusions with Fixed Suboptimal Hyperparameters
We used the same hyperparameters as in the fitting experiment to test the 1.07B model on 10B training tokens (shown in Figure 8(b)).
@@
With fixed hyperparameters, the MSE of the fitted results is $4.31\times 10^{-7}$, while using optimal hyperparameters reduces it to $2.35\times 10^{-7}$, a 45.5% improvement. This means the optimal hyperparameters are more effective for accurately capturing the model’s power-law performance.
To explore the relationship between optimal model size and compute budget, we fixed the compute budget to get the IsoFLOPs curve from Equation 40 and validated it through experiments (shown in Figure 9).
@@
Equation 38 differs from Equation 39 by an absolute error of 0.1581, or 30.26%, indicating a significant discrepancy, larger than the one observed in the fitting results under optimal hyperparameters. This discrepancy stems from the poor fit of Equation 40, particularly for lower compute budgets.

### v2

@@ 2 Setup
  • $C_{\text{token}}=\frac{3}{4}N\left(7+\frac{n_{\text{ctx}}}{d}\right)$ The compute cost per token during training, where $d$ represents the model dimension.
@@
3 Predicting Optimal Hyperparameters
@@
These analyses underscore the importance of a refined approach to hyperparameter selection in scaling law research. To address this, we propose the precise Scaling Law for video DiT (see  Appendix E for similar results on image DiT with 1 frame). In this section, we derive a power-law relationship linking hyperparameters to model size and data size, offering a theory-driven improvement over prior methods. This serves as the basis for our scaling law in Section 4.
@@ 3.1 Impact of model size on hyperparameters
Let $\mathcal{G}_{k}^{B}=\{\theta_{k},\{\xi_{k-1}^{(b)}\}_{b=1}^{B},\theta_{k-1},\{\xi_{k-2}^{(b)}\}_{b=1}^{B},\dots,\theta_{0}\}$ represent the filtration containing all historical variables at and before iteration $k$. Following [38], we assume the estimated gradient is an unbiased estimate of the true gradient, while the variance of the estimated gradient is bounded:
@@
Suppose $L(\theta_{k})$ is $L$-smooth for any iteration k  [56] and Equations 2 and 3 holds, we derive the convergence properties under the framework of mini-batch stochastic gradient descent (SGD) as in [38]. For effective updates, the learning rate $\eta$ must satisfy $\eta\leq 1/L$. Under these conditions, the convergence of $L(\theta_{k})$ is characterized as follows (proof provided in Appendix B):
@@
As the model size increases, its complexity also grows, leading to a corresponding increase in the Lipschitz constant $L$ [22]. To ensure that the upper bound on the right-hand side of Equation 4 remains manageable, both the learning rate and batch size must be adjusted with respect to $L$. Specifically, a larger $L$ necessitates a smaller learning rate $\eta$ and a larger batch size $B$ for convergence. This relationship illustrates how the scaling of model size impacts the optimal choice of hyperparameters. Unlike prior work [1], which treats model size implicitly, our analysis explicitly incorporates model size as an independent variable. This enables a more accurate estimation of optimal hyperparameters by addressing the interplay between model complexity, learning rate, and batch size in a principled manner.
@@ 3.2 Proposed $B_{\text{opt}}(N,T)$ equation
$\mathbb{D}_{\xi_{k}\sim\rho}[g_{k}]=\frac{1}{B}\Sigma_{k}$ (5)
@@ 3.4 Results of optimal hyperparameters
4 Scaling Law in Video DiT with Optimal Hyperparameters
@@ 4.1 Scaling law for optimal model size
In this section, we explore how to optimally allocate model size $N$ and training tokens $T$ within a given compute budget $C$ to maximize model performance. Building on the approach proposed in [1], we leverage the optimal hyperparameters defined in Equations 13 and 10 for our experiments. For each compute budget $[3e17,6e17,1e18,3e18,6e18]$, we conducted five experiments to ensure reliable results. As shown in Figure 6(a), we fit a parabola to each compute budget’s results to identify the minimum loss point, which corresponds to empirical optimal model size under that specific budget.
@@ 4.3.1 More parameter-efficient model sizes
We first validate our fitting results by extrapolating the compute budget to $5.85\times 10^{20}$. Under this condition, the optimal model size predicted by Equation 14 is 0.64B. Due to discrete layer selection, we conducted an experimental validation using a $14$-layer model with 719.3M parameters. According to equations Equations 10 and 13, we set the batch size to $832$ and the learning rate to $1.6\times 10^{-4}$. The resulting loss closely aligned with the prediction from Equation 15 (see Figure 5(b)), further validating our conclusions.
To validate the optimal hyperparameters yield more precise fitting results, we conducted experiments with fixed suboptimal hyperparameters to fit the relationship between optimal model size and compute budget (detailed in Appendix C). As shown in Figure 6, using fixed suboptimal hyperparameters overestimates the optimal model size, and this discrepancy grows with increasing compute budgets. At a compute budget of $1\times 10^{10}$ TFlops, the scaling law based on optimal hyperparameters predicts a model that saves $39.9\%$ of the parameters while achieving a similar validation loss. This occurs because, when computational resources are abundant relative to data, scaling the model size near the optimal size has a minimal impact on validation loss. Therefore, our scaling law precisely predict the optimal model size under optimal hyperparameters.
@@ 4.3.2 More precise validation loss predictions
The corresponding results under fixed suboptimal hyperparameters are provided in  Appendix C. We focus on the deviation in the exponent of $C$ between the predicted optimal model size and the empirical optimal model size relationships. As depicted in Figure 7, under optimal hyperparameters, the deviation results in an absolute error of $0.0148$, corresponding to a $3.57\%$ slope deviation. In contrast, as shown in Equations 38 and 37, the deviation under fixed suboptimal hyperparameters is significantly larger, with an absolute error of $0.1581$, and $30.26\%$ difference. This stark difference underscores the importance of using optimal hyperparameters to improve the precision of predicted relationships. Therefore, our scaling laws precisely predict the validation loss under optimal hyperparameters.
@@ A.3 Main Results
Based on the optimal learning rate, we fit the validation loss for any model size and training tokens ( Table 5).
@@ B.3 Stepwise Loss of Mini-Batch SGD
Following [38], the estimated gradient is unbiased, and its variance decreases inversely with the batch size $B$ (see Equations 2 and 5).
@@ Appendix C Experimental Conclusions with Fixed Suboptimal Hyperparameters
We used the same hyperparameters as in the fitting experiment to test the 1.07B model on 10B training tokens (Figure 8(b)).
@@
With fixed hyperparameters, the MSE of the fitted results is $4.31\times 10^{-7}$, while using optimal hyperparameters reduces it to $2.35\times 10^{-7}$, a $45.5\%$ improvement. This means the optimal hyperparameters are more effective for accurately capturing the model’s power-law performance.
To explore the relationship between optimal model size and compute budget, we fixed the compute budget to get the IsoFLOPs curve from Equation 39 and validated it through experiments ( Figure 9).
@@
Equation 37 differs from Equation 38 by an absolute error of $0.1581$, or $30.26\%$, indicating a significant discrepancy, larger than the one observed in the fitting results under optimal hyperparameters. This discrepancy stems from the poor fit of Equation 39, particularly for lower compute budgets.